20. The percent of fat calories that a person in America consumes each day is normally distributed with a mean of about 36 and a standard deviation of about 10. Suppose that 16 individuals are randomly chosen. Let X = average percent of fat calories.

a. For the group of 16, find the probability that the average percent of fat calories consumed is more than 5  ___________Answer

b. Find the first quartile for the average percent of fat calories. Round answer to two decimal places (i.e. 1.23) __________Answer
Answer #1

mean, mu = 36
std. dev. sigma = 10

Here population std. dev. is given, hence we can use z-values to calculate the probability as per central limit theorem.

z = (xbar - mu)/(sigma/sqrt(n))

a)
P(X > 5)
= P(z > (5 - 36)/(10/sqrt(16)))
= P(z > -12.4)
= 1

This value of z is very extreme to the left, hence probability to its right is 1

b)
lets find the value of z which represents the first quartile in the normal curve
P(z < -0.67) = 0.25

z = -0.67

x = 36 -0.67*10/sqrt(16)
x = 34.33